OpenStudy (anonymous):

how do i find the slope intercept form of 3x-2y=-16?

OpenStudy (amistre64):

solve for y

OpenStudy (amistre64):

3x-2y = -16 ; -3x -2y = -3x - 16 ; /-2 y = -3x/-2 -16/-2 y = (3/2)x +8

OpenStudy (anonymous):

what is the -3x for?

OpenStudy (anonymous):

is y=(3/2)x+8 the answer?

OpenStudy (amistre64):

when we have an equal sign in there that means that both saides are balanced; they weigh the same

OpenStudy (amistre64):

we have to keep them balanced so whatever we do to one side we have to do to the other....right

OpenStudy (amistre64):

how do we get y all by its little lonesome? first we have to -3x from it right?

OpenStudy (anonymous):

can you walk me through the rest of the problem?

OpenStudy (amistre64):

3x-3x -2y = -16-3x 0 -2y = -3x-16 -2y = -3x-16 do you agree?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

now we have a -2 stuck the the 'y'...and its stuck there by multiplication....we remove it by dividing. (-2/-2)y = (-3x/-2) -(16/-2) 1y = (3/2)x + 8 y = (3/2)x+8 ; do you agree?

OpenStudy (amistre64):

by doing albegraic techniques we simply change the way the problem looks; not the value of it...jus tthe shape of it

OpenStudy (anonymous):

yes! so y equals (3/2)x+8?

OpenStudy (amistre64):

unless you came to another conclsion without me...then yes, the final answer is the final answer :)

OpenStudy (anonymous):

thank you!